# RoCo-Sim: Enhancing Roadside Collaborative Perception through Foreground Simulation

### v1

@@ Abstract
Roadside Collaborative Perception refers to a system where multiple roadside units collaborate to pool their perceptual data, assisting vehicles in enhancing their environmental awareness. Existing roadside perception methods concentrate on model design but overlook data issues like calibration errors, sparse information, and multi-view consistency, leading to poor performance on recent published datasets. To significantly enhance roadside collaborative perception and address critical data issues, we present the first simulation framework RoCo-Sim for road-side collaborative perception. RoCo-Sim is capable of generating diverse, multi-view consistent simulated roadside data through dynamic foreground editing and full-scene style transfer of a single image. RoCo-Sim consists of four components: (1) Camera Extrinsic Optimization ensures accurate 3D to 2D projection for roadside cameras; (2) A novel Multi-View Occlusion-Aware Sampler (MOAS) determines the placement of diverse digital assets within 3D space; (3) DepthSAM innovatively models foreground-background relationships from single-frame fixed-view images, ensuring multi-view consistency of foreground; and (4) Scalable Post-Processing Toolkit generates more realistic and enriched scenes through style transfer and other enhancements. RoCo-Sim significantly improves roadside 3D object detection, outperforming SOTA methods by 83.74% on Rcooper-Intersection and 83.12% on TUMTraf-V2X for AP70. RoCo-Sim fills a critical gap in roadside perception simulation. Code and pre-trained models will be released soon: https://github.com/duyuwen-duen/RoCo-Sim
@@ 1 Introduction
To address these challenges, we introduce RoCo-Sim, the first simulation framework designed for roadside collaborative perception. RoCo-Sim is capable of generating diverse, multi-view consistent simulated roadside data through dynamic foreground editing and full-scene style transfer of a single image, as shown in Fig. 2. Since the editing is performed in 3D space, our modifications made from one viewpoint can automatically propagate to other viewpoints, ensuring spatial consistency across multiple perspectives. The key design rationale behind RoCo-Sim is to leverage a rich 3D asset library to establish a 3D-to-2D mapping and, with the aid of graphical tools [3], seamlessly render foreground objects onto real 2D backgrounds. RoCo-Sim encompasses 4 key components: (i) Camera Extrinsics Optimization for accurate camera pose modeling; (ii) A novel Multi-View Occlusion-Aware Sampler (MOAS), which determines the occlusion-aware placement of digital assets within the 3D space, ensuring that assets are dynamically positioned in a physically plausible manner; (iii) DepthSAM, which innovatively models the foreground-background relationships in each viewpoint, enabling occlusion-aware insertion of 3D assets while maintaining multi-view consistency; and (iv) Scalable Post-Processing Toolkit for seamless integration through style transfer and other enhancements.
Extensive experiments are conducted on two real-world roadside datasets, RCooper [8] and TUMTraf-V2X [47], to evaluate RoCo-Sim. Our findings indicate that: i) Camera Extrinsics Optimization greatly improves model performance. Performance on RCooper improves by 62.55% for AP70. ii) The performance improvement for perception becomes more significant as the amount of simulation data and the number of simulated vehicles per image increase. iii) RoCo-Sim significantly enhances perception performance. It enhances previous SOTA performance by 83.74% on Rcooper-Intersection and 83.12% on TUMTraf-V2X for AP70. To sum up, our contributions are:
@@ 3.3 DepthSAM
We denotes $D_{\text{rel}}\in\mathbb{R}^{H\times W}$ represents the relative depth predicted by DepthAnything [41] and $Z\in\mathbb{R}^{H\times W}$ represents the projected depth map. The projected depth is obtained from a 3D point cloud $P_{3d}\in\mathbb{R}^{N\times 3}$ , where N is the number of points. The point cloud is projected into the camera coordinate system using $P_{\text{cam}}=\mathcal{M}_{\text{l2c}}\cdot P_{3d}$. The depth value $Z(u,v)$ satisfying $Z(u,v)\cdot(u,v,1)^{T}=\mathcal{K}\cdot P_{cam}$ for pixel $(u,v)$. We solve for the calibration parameters a and b using the following optimization problem:
@@ 4 Experiment
Table 2: RoCo-Sim greatly enhances detection performance, even surpassing those achieved by the optimized model BEVSpread.
@@ 4.5 Ablation studies
Camera extrinsic optimization algorithm. From Tab. 4, we find that calibrating the camera extrinsics can significantly enhance model performance. For instance, in the case of 117, as shown in Fig. 2, the extrinsics are quite inaccurate, resulting in a 62.55% improvement in performance for AP70 after calibration. This is because camera extrinsics calibration aligns the images and labels, making it easier for the model to learn useful information from them.
@@ 5 Conclusion
This work unleashes the power of simulation data to enhance roadside collaborative perception, demonstrating that simulation data, rather than model architecture, is the true winner in performance competition. The proposed RoCo-Sim is the first simulation framework specifically designed for roadside collaborative perception. RoCo-Sim generates a large scale of multi-view consistent simulation data from sparse fixed viewpoints, supports both foreground editing and full-scene style transformation, and can be rapidly deployed to any new road environment. We hope this work will significantly accelerate the practical development of roadside collaborative perception, ultimately enhancing driving safety.

### v3

@@ Abstract
Roadside Collaborative Perception refers to a system where multiple roadside units collaborate to pool their perceptual data, assisting vehicles in enhancing their environmental awareness. Existing roadside perception methods concentrate on model design but overlook data issues like calibration errors, sparse information, and multi-view consistency, leading to poor performance on recent published datasets. To significantly enhance roadside collaborative perception and address critical data issues, we present the first simulation framework RoCo-Sim for road-side collaborative perception. RoCo-Sim is capable of generating diverse, multi-view consistent simulated roadside data through dynamic foreground editing and full-scene style transfer of a single image. RoCo-Sim consists of four components: (1) Camera Extrinsic Optimizer ensures accurate 3D to 2D projection for roadside cameras; (2) A novel Multi-View Occlusion-Aware Sampler (MOAS) determines the placement of diverse digital assets within 3D space; (3) DepthSAM innovatively models foreground-background relationships from single-frame fixed-view images, ensuring multi-view consistency of foreground; and (4) Scalable Post-Processing Toolkit generates more realistic and enriched scenes through style transfer and other enhancements. RoCo-Sim significantly improves roadside 3D object detection, outperforming SOTA methods by 83.74% on Rcooper-Intersection and 83.12% on TUMTraf-V2X for AP70. RoCo-Sim fills a critical gap in roadside perception simulation. Code will be released soon: https://github.com/duyuwen-duen/RoCo-Sim
@@ 1 Introduction
To address these challenges, we introduce RoCo-Sim, the first simulation framework designed for roadside collaborative perception. RoCo-Sim is capable of generating diverse, multi-view consistent simulated roadside data through dynamic foreground editing and full-scene style transfer of a single image, as shown in Fig. 2. Since the editing is performed in 3D space, our modifications made from one viewpoint can automatically propagate to other viewpoints, ensuring spatial consistency across multiple perspectives. The key design rationale behind RoCo-Sim is to leverage a rich 3D asset library to establish a 3D-to-2D mapping and, with the aid of graphical tools[3], seamlessly render foreground objects onto real 2D backgrounds. RoCo-Sim encompasses 4 key components: (i) Camera Extrinsics Optimizer for accurate camera pose modeling; (ii) A novel Multi-View Occlusion-Aware Sampler (MOAS), which determines the occlusion-aware placement of digital assets within the 3D space, ensuring that assets are dynamically positioned in a physically plausible manner; (iii) DepthSAM, which innovatively models the foreground-background relationships in each viewpoint, enabling occlusion-aware insertion of 3D assets while maintaining multi-view consistency; and (iv) Scalable Post-Processing Toolkit for seamless integration through style transfer and other enhancements.
Extensive experiments are conducted on two real-world roadside datasets, RCooper [8] and TUMTraf-V2X [47], to evaluate RoCo-Sim. Our findings indicate that: i) Camera Extrinsics Optimizer greatly improves model performance. Performance on RCooper improves by 62.55% for AP70. ii) The performance improvement for perception becomes more significant as the amount of simulation data and the number of simulated vehicles per image increase. iii) RoCo-Sim significantly enhances perception performance. It enhances previous SOTA performance by 83.74% on Rcooper-Intersection and 83.12% on TUMTraf-V2X for AP70. To sum up, our contributions are:
@@ 3.3 DepthSAM
[…]
@@ 4 Experiment
Table 2: RoCo-Sim greatly enhances detection performance, with the improvements brought by simulation even surpassing those achieved by the optimized model BEVSpread.
@@ 4.5 Ablation studies
Camera Extrinsic Optimizer. From Tab. 4, we find that calibrating the camera extrinsics can significantly enhance model performance. For instance, in the case of 117, as shown in Fig. 2, the extrinsics are quite inaccurate, resulting in a 62.55% improvement in performance for AP70 after calibration. This is because camera extrinsics calibration aligns the images and labels, making it easier for the model to learn useful information from them.
@@ 5 Conclusion
This work unleashes the power of simulation data to enhance roadside collaborative perception, demonstrating that simulation data, rather than model architecture, is the true winner in performance competition. The proposed RoCo-Sim is the first simulation framework specifically designed for roadside collaborative perception. RoCo-Sim generates a large scale of multi-view consistent simulation data from sparse fixed viewpoints, supports both foreground editing and full-scene style transformation, and can be rapidly deployed to any new road environment. We hope this work accelerates practical roadside collaborative perception development, ultimately enhancing driving safety.